Required information [The following information applies to the questions displayed below.) Hills Companys June 30 bank state2. Prepare any journal entries that should be made as a result of the bank reconciliation. (If no entry is required for a tra3. What is the balance in the Cash account after the reconciliation entries? Cash Balance4. In addition to the balance in its bank account, Hills Company also has $260 of petty cash on hand. This amount is recorded
Answer #1
Bank service charges 30
Cash 30

3.

Cash balance = 6560 - 30 = 6530

4.

Totsl amount of cash = 6530 + 260 = 6790